RHIC Startup: Gold injection has been a problem since the flip of the PHENIX magnet polarity.
Power supply problems with b02-qgt required a new jump card to fix Blue losses at transition. An access was made to Alcove 3 to replace the card.
Last night's fills were interrupted by two RF problems. Blue Storage cavity 3 (BS3) had trouble turning on because of low temperature interlock. Yellow Accelerating Cavity 1 (YA1) had a bad relay in the PASS indicator. An access was required to turn on the Ignitron of BS3. When they were brought back on, beam could be filled and stored.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |